# On the CP Nature of the ‘95 GeV’ Anomalies

### v2

@@ 4 Results
[…]
Now, we show how to uncover the CP properties of the scalar particle at the LHC. For a realistic representation of signal and background events, we have used the present CMS data of the $\tau^{+}\tau^{-}$ excess CMS:2018cyk ; CMS:2023yay . The data is taken from HEPDATAMaguire:2017ypu ; hepdata.128147 . For our analysis, we have selected the events for which the invariant mass $m_{\tau^{+}\tau^{-}}$ lies within $60-120$ GeV and we have selected the events with $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV obtained by the vectorial $p_{T}$ sum of the visible decay products of tau lepton. This helps to minimise the background events significantly, as evident from Fig. 9 in ref CMS:2022goy .
@@
To demonstrate the CP nature of the scalar, we plot the $\phi_{CP}$ for both signal and background events. The distribution peaks at a particular value of $\phi_{CP}$ for a given CP-admixture, whereas the background remains primarily flat. This enables estimating the CP nature from the $\phi_{CP}$ distribution alone. We found that with an integrated luminosity of 138 $fb^{-1}$, the total event distribution remains within statistical uncertainty of the background ($\sqrt{N_{B}}$). We considered high luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw , and appropriately rescaled the existing data to an integrated luminosity of 3000 $fb^{-1}$. The $\phi_{CP}$ distribution for signal and background events for HL-LHC is shown in Fig. 2, where background events are shown in yellow with statistical uncertainty bars, and the signal events are shown in green. As anticipated, the background events exhibit a near-uniform distribution in all the cases. The top panel depicts the distribution for CP-Even (top-left) and CP-odd (top-right) scalar, whereas the bottom panels depict maximally CP-Mixed states. From all the figures, it is evident that for a given CP nature of the resonance, the cumulative number of signal and background events rise beyond the background uncertainty, signalling the CP property of the scalar.
$\alpha$ $\phi_{CP}$ Window $\mathcal{S}_{0}$ $\mathcal{S}_{\sqrt{b}}$ $\mathcal{S}_{1\%}$ $\mathcal{S}_{10\%}$
00 $120^{\circ}-240^{\circ}$ $3.10$ $2.18$ $2.01$ $0.26$
$\pi/2$ $0^{\circ}-80^{\circ}~{}~{}\&~{}~{}280^{\circ}-360^{\circ}$ $3.51$ $2.47$ $2.09$ $0.26$
$\pi/4$ $200^{\circ}-320^{\circ}$ $3.13$ $2.2$ $2.03$ $0.27$
-$\pi/4$ $40^{\circ}-160^{\circ}$ $3.12$ $2.2$ $2.03$ $0.27$
Table 1: Significance of different CP eigenstates with various background uncertainties at HL-LHC.
In order to quantify the result, we have chosen a window in $\phi_{CP}$ distribution and estimated a signal significance using the following formula Cowan:2010js ,
$\mathcal{S}_{\sigma_{b}}=\sqrt{2}\Big{\{}(N_{S}+N_{B})\ln\Big{[}\frac{(N_{S}+N_{B})(N_{B}+\sigma_{b}^{2})}{N_{B}^{2}+(N_{S}+N_{B})\sigma_{b}^{2}}\Big{]}-\frac{N_{B}^{2}}{\sigma_{b}^{2}}\ln\Big{[}1+\frac{\sigma^{2}_{B}N_{S}}{N_{B}(N_{B}+\sigma_{b}^{2})}\Big{]}\Big{\}}^{1/2},$ (17)
where $N_{S(B)}$ denotes signal(background) events and $\sigma_{b}$ denotes the background uncertainty. In Tab. 1 we have tabulated the significance for an optimal $\phi_{CP}$ window with various background uncertainty to achieve a robust estimation of the Significance. As expected, vanishing BG uncertainty provides a significance of more than $3~{}\sigma$ at HL-LHC. As uncertainty increases, Significance reduces but remains large enough if we consider the background uncertainty solely statistical ($\sqrt{N_{B}}$) or $1\%$. If $\sigma_{b}$ becomes large ($\sim 10\%$), the number of signal events need to be higher to achieve any significance.
@@ 5 Conclusions
In summary, under the assumption that several anomalies seen in the invariant mass of $\gamma\gamma$, $\tau^{+}\tau^{-}$ (from LHC data) and $b\bar{b}$ (from LEP data) events are due to a single spin-0 object with mass 95 GeV having Yukawa couplings to third generation fermions, we have performed a detector level analysis aimed at extracting the CP properties of such an object. We have done so limitedly to the $\tau^{+}\tau^{-}$ final state, as it offers the possibility of accessing the corresponding CP quantum numbers whichever the $\tau$ decay, whether leptonic or hadronic, through the relative angle between the two decay planes defined through the visible particles emerging from the $\tau$ lepton decays. Our approach has been based on a simplified model, which is characterized by the relevant Yukawa interactions being defined in terms of correcting factors to the corresponding SM vertex strengths while allowing for both a scalar and pseudoscalar interaction as well as any superposition thereof.
We have found that, under the assumption that the $\gamma\gamma$ and $\tau^{+}\tau^{-}$ anomalies will persist throughout the LHC era, the HL-LHC, in standard energy and luminosity configurations, will offer the chance to extract the CP properties of such a new spin-0 resonance. Once the latter is combined with cross-section and BR information, it would then be possible to map the results obtained in our model-independent parameterization into well-defined models like, e.g., a 2HDM of whichever Type, so long that it can host a suitable 95 GeV state in its Higgs particle spectrum, as it has indeed been proven to be the case in previous literature.
@@ Acknowledgements
T.M. and P.S. would like to thank Artur Gottmann and Ravindra Kumar Verma for some useful discussions. T.M. is supported by the BITS Pilani Grant NFSG/PIL/2023/P3801. S.M. is supported in part through the NExT Institute and STFC Consolidated Grant No. ST/L000296/1.

### v3

@@ 4 Results
Figure 2: Signal and background distribution of our CP-sensitive observable for combining all the hadronic (left panels) and semi-leptonic (right panels) modes with an integrated luminosity of 3000 fb-1. The top panel shows the distribution for a CP-even 95 GeV scalar, while the middle panel shows the same for a maximally CP-mixed state, and the CP-odd distribution is shown in the bottom panel.
Figure 3: Signal and theoretical distribution fo the CP observable $\phi_{\rm CP}$ scaled to HL-LHC from the observed data at 139 fb-1. The top panels show the distribution for CP-even (top-left) and maximally CP-mixed (top-right) scalar, and the bottom panels show the distribution for a CP-odd state. The hadronic IP method provides the best reconstruction and closely resembles the theoretical distribution, whereas the other methods show the exact nature of the distribution with subdued extrema. If $\alpha$ becomes $\pi/4$, then the peak of the distribution will be at $\phi_{\rm CP}=\pi/2$.
@@
Now, we show how to access the CP properties of the spin-0 with particle with mass around 95 GeV at the LHC. For a realistic representation of signal and background events, we have used the present CMS data of the $\tau^{+}\tau^{-}$ excess CMS:2018cyk ; CMS:2023yay . The data is taken from HEPDATA and for our analysis, we have selected the events for which the invariant mass $m_{\tau^{+}\tau^{-}}$ lies within 6012060-120 GeV with $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV. To demonstrate the CP nature of the spin-0 object, we obtain the $\phi_{\rm CP}$ distribution for both signal and background events. We found that, with an integrated luminosity of 138 fb-1, the total event distribution remains within statistical uncertainty of the background ($\sqrt{N_{B}}$) and that the situation is not considerably better at the end of Run 3. Hence, we considered the HL-LHC Gianotti:2002xx ; Apollinari:2015wtw and appropriately rescaled the existing data to an integrated luminosity of 3000 fb-1. The $\phi_{\rm CP}$ distribution for signal and background events for the HL-LHC is shown in Fig. 2, where background events are shown in yellow, whereas the signal events are shown in green. The combined hadronic events are shown in the left panel, whereas the right panel depicts semi-leptonic events after combining all the modes. The top, middle and bottom panels in Fig. 2 depicts the distribution for a CP-even, maximally CP-mixed and CP-odd state, respectively. As anticipated, the background events exhibit a near-uniform distribution in all cases. As theoretically expected, the signal distribution peaks at $\phi_{\rm CP}=\pi(0\textrm{ and }2\pi)$ for a CP-even(odd) spin-0 state, whereas for the CP-mixed one, the peak is at $3\pi/2$. From all the figures, it is evident that, for a given CP nature of the resonance, the cumulative number of signal events rises beyond the background uncertainty, thus signalling the CP properties of the new particle.
Figure 4: Plot of $\Delta\chi^{2}$ as a function of fitting angle $\alpha_{fit}$ assuming that the 95 GeV scalar to be CP-even(red), maximal CP-mixed(blue) and CP-odd(green). The solid(dashed) line corresponds to background systematics assumed to be 0.5%(1%). The black horizontal line at $\Delta\chi^{2}=13.36$ shows the 90% Confidence Level (CL) interval.
To quantify our results, we use the $\chi^{2}$ defined as
$\chi^{2}(\alpha)=\sum_{\rm Modes}\quad\sum_{i\in\rm{Bins}}\dfrac{\left(S_{i}^{\alpha_{H}}-\frac{n_{S}}{\Gamma}\frac{d\Gamma}{d\phi_{\rm CP}}(\alpha)\right)^{2}}{(\delta S_{i})^{2}+(\delta_{sys}^{2})},$ (17)
where $\alpha_{H}$ is the assumed hypothesis (i.e., the CP nature) for the 95 GeV state and $S_{i}^{\alpha_{H}}$ denotes the signal events in the $i^{th}$ bin. The input parameter $\alpha$ provides the theoretical estimation of the CP observable $\phi_{CP}$ normalized to the total number of signal events, $n_{S}$. For each bin, the variance includes the statistical component defined as $\delta S_{i}=\sqrt{S_{i}}$ and an uncorrelated systematic uncertainty corresponding to 0.5% or 1% of the expected background in that bin. Finally, the summation over modes includes all the semi-leptonic and hadronic $\tau^{+}\tau^{-}$ divided into IP, $\rho$ and combined IP-$\rho$ modes. The normalized binned distribution of signal events $(S_{i}^{\alpha_{H}})$ as well as the theoretical distribution $\left(\dfrac{n_{S}}{\Gamma}\dfrac{d\Gamma}{d\phi_{\rm CP}}(\alpha)\right)$ for various channels is shown in Fig. 3 for a CP-even (top-left), maximally mixed-CP state (top-right) as well as CP-odd (bottom). As expected, the hadronic IP method yields the most accurate distribution, closely matching the theoretical distribution, while the other methods capture the overall shape, but with maxima and minima systematically suppressed relative to the theoretical distribution.
In Fig. 4, we present our main result: the distribution of $\Delta\chi^{2}$ as a function of $\alpha_{\text{fit}}$. The red, blue and green lines correspond to the assumed hypotheses that the 95 GeV scalar is CP-even, maximally CP-mixed, and CP-odd, respectively. The solid (dashed) curves represent $\delta_{\text{sys}}=0.5\%\ (1\%)$ of the total background, which enters the $\chi^{2}$ calculation as defined in Eq. (17). The signal distribution is binned with $N_{\text{bin}}=9$ and the $90\%$ CL interval corresponds to $\Delta\chi^{2}=13.36$, shown as a black horizontal line in Fig. 4. We find that, at the HL-LHC, it will be possible to ascertain the CP nature of the 95 GeV scalar within $\pm 0.27\ (0.47)$ radians of the true hypothesis at $90\%$ CL, depending on the assumed background systematics.
@@ 5 Conclusions
In summary, under the assumption that several anomalies seen in the invariant mass of $\gamma\gamma$, $\tau^{+}\tau^{-}$ (from LHC data) and $b\bar{b}$ (from LEP data) events are due to a single spin-0 object with mass approximately 95 GeV having Yukawa couplings to third generation fermions, we have performed a detector level analysis aimed at extracting the CP properties of such an object. We have done so limitedly to the $\tau^{+}\tau^{-}$ final state, as it offers the possibility of accessing the corresponding CP quantum numbers, whichever the $\tau$ decay, whether leptonic or hadronic, through the relative angle between the two decay planes defined through the visible particles emerging from the $\tau$ lepton decays. Our approach has been based on a simplified model, which is characterized by the relevant Yukawa interactions being defined in terms of correcting factors to the corresponding SM vertex strengths while allowing for both a scalar and pseudoscalar interaction as well as any superposition thereof.
We have found that, under the assumption that the $\gamma\gamma$ and $\tau^{+}\tau^{-}$ anomalies will persist throughout the LHC era, the HL-LHC, in standard energy and luminosity configurations, will offer the chance to extract the CP properties of such a new spin-0 resonance. Once the latter is combined with cross section and BR information, it would then be possible to map the results obtained in our model-independent parameterization into well-defined models like, e.g., a 2HDM of whichever Type, so long that the latter can host a suitable 95 GeV state in its Higgs particle spectrum.
@@ Acknowledgements
T.M. and P.S. would like to thank Artur Gottmann and Ravindra Kumar Verma for some useful discussions. T.M. is supported by the BITS Pilani Grant NFSG/PIL/2023/P3801. S.M. is supported in part through the NExT Institute and STFC Consolidated Grant No. ST/L000296/1. P.S. is supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) of Sogang University (RS-2020-NR049598). P.S. would also like to thank the KIAS Center for Advanced Computation for providing computing resources.